### v1

# Multiway Alignment of Political Attitudes

Abstract
The related concepts of partisan belief systems, issue alignment, and partisan sorting are central to our understanding of politics. These phenomena have been studied using measures of alignment between pairs of topics, or how much individuals’ attitudes toward a topic reveal about their attitudes toward another topic. We introduce a higher-order measure that extends the assessment of alignment beyond pairs of topics by quantifying the amount of information individuals’ opinions on one topic reveal about a set of topics simultaneously. Our multiway alignment measure indicates how much individuals’ opinions on multiple topics align into a single ideological divide. Applying this approach to legislative voting behavior reveals that parliamentary systems typically exhibit similar multiway alignment characteristics, but can change in response to shifting intergroup dynamics. In American National Election Studies surveys, our approach reveals a growing significance of party identification together with a consistent rise in multiway alignment over time. Similarly, the growing multiway alignment among topical issues in Finnish online discussions suggests a trend towards a more ideologically driven political landscape. Our case studies demonstrate that the multiway alignment measure is a versatile tool for understanding societal polarization and partisan belief systems across diverse domains.
1 Introduction
The concept of belief systems, which encompass configurations of ideas and attitudes constrained by some form of interdependence (Converse, 2006), has long been central to the study of politics. Extensive evidence suggests that the emergence of political belief systems influences political affiliations, attitudes, judgments, behaviors, and the perception of facts (Converse, 2006; Van Bavel and Pereira, 2018). Strong political belief systems may contribute to the formation of entrenched partisan identities, which threatens the democratic process, leading to conflicts and hindering the ability of individuals to understand outgroup views (Hetherington, 2009).
Constraint between attitudes is often understood as the extent to which an individual’s position toward one issue predicts their position toward another (Converse, 2006). Similarly, partisan sorting manifests as individuals taking on the issue positions of the parties they identify with (Baldassarri and Gelman, 2008). The emergence of partisan belief systems, then, can be investigated by considering how attitudes toward issues are aligned, through some form of statistical association, with attitudes toward other issues (i.e., issue alignment) or toward parties (i.e., partisan alignment or issue partisanship). Previous research on partisan sorting, issue alignment, and political polarization has done so using pairwise metrics, such as correlation (Baldassarri and Gelman, 2008; Kozlowski and Murphy, 2021; Bougher, 2017; Hetherington, 2009; Lüders et al., 2024) and mutual information (Chen et al., 2021).
However, we have no reason to believe constraints only exist at the pairwise level – a belief system contains a collection of issue pairs, which are unlikely to be independent of one another. While existing pairwise approaches intuitively capture the concept of constraints in belief systems, and can even be used to study collections of issue pairs (e.g., Lüders et al., 2024; Fishman and Davis, 2022; Chen et al., 2021), they are unable to describe alignment jointly across multiple topics (i.e., issues and parties), missing the potential for there to be multidimensional constraints. Exclusively focusing on pairwise associations may fail to capture the richer structure of higher-order interdependencies that exists among sets of political topics – such as how voters align themselves on multiple issues simultaneously or how legislators position themselves in a multidimensional policy space – which, as we show below, is not simply the aggregation of pairwise constraints.
To better understand how belief systems shape political attitudes, we propose the careful consideration of how topics can be jointly aligned, and we develop a measure of multiway alignment to study this phenomenon. By pursuing multiway alignment, we are emphasizing that real-world systems extend beyond pairwise interactions (Battiston et al., 2020; Bougher, 2017; Levin et al., 2021), a notion that intuitively flows from our existing understanding of politics. An example of this can be found in the liberal versus conservative divide in U.S. politics (Bougher, 2017). Numerous studies on partisan sorting demonstrate that individuals who identify as liberal or conservative are more likely to align themselves with the party that offers policy bundles consistent with their ideological leanings (Bafumi and Shapiro, 2009; Baldassarri and Gelman, 2008; Bougher, 2017). More recently, what had primarily been a process of partisan alignment has also extended to entrenched ideological or issue alignment (Kozlowski and Murphy, 2021). This phenomenon is not limited to the U.S., as similar trends towards polarization have also been observed in European countries (Westwood et al., 2018; Flores et al., 2022). These observations suggest that the study of polarization and partisan identities should focus on cleavages arising from the alignment of multiple issues.
When examining ideological cleavages across multiple topics and multidimensional constraints, researchers can use multivariate measures that capture global characteristics of joint distributions (e.g., Rosas et al., 2019), such as the extent of information shared among variables and the level of interdependence, through the entropies of the stochastic variables. We document these in E. Some of these measures, however, present applicational challenges in that they do not have directly correspondent bivariate analyses. Moreover, the use of unnormalized measures complicates the interpretability of the scores and comparisons across different datasets.
We derive a measure of multiway alignment based on entropy and mutual information to capture higher-order interdependencies. Our proposed quantification of multiway alignment generalizes the method in Chen et al. (2021), allowing to quantify alignment across any number of topics. While pairwise alignment in Chen et al. (2021) is quantified by the amount of information that knowing someone’s opinion on one topic provides on their opinion on another topic, we measure the amount of information that knowing someone’s opinion on one topic provides on their opinion on the remaining $n-1$ topics. Additionally, we ensure the significance of our proposed score by comparing it with the amount of alignment that can be attributed to chance.
Our approach to multiway alignment analysis facilitates understanding the interplay between various divides and beliefs. We illustrate this using multiple cases that vary by their political and geographical context. Our analysis of data from the American National Election Studies (ANES; 2021) reveals the increasingly substantial influence of party identification in shaping opinion dynamics among the U.S. public. In Finnish online political discourse, we observe a pronounced increase in higher-order alignment across issues, signaling a transformation towards a more ideologically driven political environment. Our examination of the Finnish Parliament before and during the COVID pandemic highlights a shift toward considerably lower levels of higher-order alignment in response to the challenges posed by the crisis, a well-recognized phenomenon that pairwise analysis alone does not capture. Finally, we observe elite polarization from voting behavior in the U.S. House, where a multitude of policy issues manifests high alignment. To facilitate further exploration of our findings and application to other systems, we provide the Python package multiway_alignment (H), which lets researchers apply our multiway alignment methodology.
The paper proceeds as follows. Section 2 establishes the theoretical framework that extends the concept of alignment beyond the pairwise to effectively model higher-order interdependencies. Section 3 details our approach to quantifying multiway alignment. In Section 4, we demonstrate the broad applicability of our measure to a variety of social and political systems and we discuss our findings. Finally, Section 5 summarizes our contributions and outlines future research directions.
2 Importance of Multiway Alignment
Figure 1: Illustration of consensus partitions based on individuals’ preferences toward three issues (tie color $T_{1}$, shirt color $T_{2}$, and pants color $T_{3}$). Panel (a) illustrates a system with no alignment, where the opinion groups found within each issue become fragmented into smaller consensus groups as additional issues are considered. Panel (b) illustrates a system with perfectly aligned positions on the three issues. In this case, the resulting consensus partition exactly matches the opinion groups for each issue.
We propose a multiway alignment measure that gauges the extent to which individuals’ attitudes toward different topics collapse into a single dimension. This concept goes beyond examining alignment over pairs of topics and instead investigates the alignment of opinions over an array of topics and the extent to which consensus among individuals transcends the specific subjects being discussed. To do so, we define the idea of consensus partition to capture the multidimensional nature of social systems. Specifically, a consensus partition refers to a division of individuals into groups (i.e., a consensus group) based on how they share opinions across multiple topics. As such, these partitions comprise groups of individuals who hold the same constellation of positions across a set of topics. We further formalize this concept in Section 3.
Consensus partitions matter to our understanding of alignment across multiple topics in that they capture the extent to which ideological or partisan cleavages sort individuals into groups that have overlapping preferences, which depends on the extent to which these cleavages are either aligned or cross-cutting. Specifically, when individuals can take one of two positions on any political topic, a system with $n$ topics means there are $2^{n}$ possible constellations of positions (i.e., consensus groups). When cleavages are highly aligned, individuals will sort into a consensus partition such that only few of these consensus groups actually meaningfully exist (i.e., most groups are very small or nonexistent). At the extreme where cleavages are perfectly aligned, there will only be two consensus groups that have no shared attitudes between them. On the other hand, when cleavages are cross cutting, there will be many relatively balanced consensus groups in society, and even relatively like-minded individuals will rarely agree on everything, and in turn will not disagree with their outgroup on everything. In this respect, cross-cutting cleavages create a more complex identity space, as the diversity of stances results in a tapestry of identities that defy simple categorization (e.g., someone might be economically conservative but socially liberal, or support environmental policies while holding different views on immigration).
We illustrate this concept more explicitly in Figure 1 with two extreme cases involving eight individuals, and three issues with binary choices (i.e., green or purple ties, shirts, and pants). The absence of alignment between the three issues (Figure 1a) yields the least aligned consensus partition of maximally balanced consensus groups ($82^{3}=$ 8 groups with one person in each group). In this case, even when the issues themselves are divisive, the impact of exclusive identities is mitigated since opinion groups toward one issue are not related to opinion groups toward the other issues. Individuals can find common ground despite disagreement on specific issues – for all individuals in Figure 1a, there is only one person with whom they have no shared preference at all.
On the other extreme, Figure 1b shows an instance of consistent alignment of opinion groups across the three issues, indicating high interdependence among the issues. This means the individuals in the two different consensus groups hold conflicting views on all three issues. Their consistent opposition implies that the two groups’ beliefs are structured using a similar rationale, albeit applied in opposite directions. The implication here is that, despite the even split of individuals within each issue in both examples, the perfect alignment of issues in the latter case yields large and entrenched consensus groups of individuals who always agree with their ingroup and never with the outgroup.
In general, each political topic either introduces potential for cross-cutting cleavages or confirms (i.e., aligns with) the existing consensus partition. To illustrate this, consider the addition of a fourth issue, e.g., preference toward shoes, which again has four green positions and four purple positions. If shoe preferences align with existing preferences for pants, shirts, and ties, all individuals who prefer green pants, green shirts, and green ties will also agree on the same color of shoes. This alignment would strengthen cohesive preferences across the wider array of issues without altering the existing consensus groups – there would still be two groups even though there are $162^{4}=$ 16 possible ones. Conversely, if shoe preferences do not align with other issues under consideration, individuals who prefer green pants, green shirts, and green ties may differ in their opinions on shoe color. Some may opt for green shoes while others choose purple shoes. This lack of alignment would fragment the previously established consensus partition, resulting in additional, balanced consensus groups.
2.1 Pairwise Alignment Misses Higher-Order Patterns
Figure 2: An example where 2-way alignment measures cannot capture multiway alignment. Panel (a) and Panel (b) show two systems of ten individuals exhibiting the exact same set of pairwise associations but different multiway alignment over three topical issues ($T_{1},T_{2},T_{3}$). The squared cells represent individuals’ preference on each issue ("agree" in purple, and "disagree" in green). Issue $T_{1}$ and issue $T_{2}$ are the same in Panel (a) and in Panel (b), whereas $T_{3}$ and $T^{\prime}_{3}$ generate different opinion partitions. The consensus partitions $C(T_{2},T_{3})$ and $C(T_{2},T^{\prime}_{3})$ show that the amount of pairwise alignment between issue $T_{2}$ and $T_{3}$ and between $T_{2}$ and $T^{\prime}_{3}$ is exactly the same, since both pairs generate the same number of consensus groups, and individuals are partitioned with the same relative frequencies. The same holds for $T_{1}$ and $T_{3}$ in Panel (a) and for $T_{1}$ and $T^{\prime}_{3}$ in Panel (b). Yet, $C(T_{1},T_{2},T_{3})$ in Panel (a) and $C(T_{1},T_{2},T^{\prime}_{3})$ in Panel (b) demonstrate that the pairwise alignments do not uniquely determine multiway alignment: $C(T_{1},T_{2},T_{3})$ in Panel (a) and $C(T_{1},T_{2},T^{\prime}_{3})$ in Panel (b) describe two different amounts of 3-way alignment, lower in Panel (a) and higher in Panel (b).
As noted above, while one could use pairwise alignment measures between all pairs of topics to estimate the amount to which all of them are aligned – for the example in Figure 1, this would amount to computing three values, one between each of the three pairs – it would be insufficient for inferring the existence of higher-order constraints. Measures such as correlation or mutual information can describe the extent of pairwise alignment, but they ignore the potential dependencies that may exist among three or more topics simultaneously. Therefore, summarizing a set of pairwise scores, for example by summing or averaging, to describe a higher-order constraint can lead to an inaccurate representation of the higher-order constraints.
We illustrate two potential shortcomings of pairwise measures using the examples in Figure 2. First, the level of alignment in a system at the pairwise level does not necessarily reflect the level of alignment at the triad and higher. Consider Figure 2a, which illustrates three issues ($T_{1}$, $T_{2}$, and $T_{3}$) with binary choices (represented by color, i.e., green or purple). The consensus partitions involving only two issues (i.e., $C(T_{1},T_{2})$, $C(T_{1},T_{3})$, and $C(T_{2},T_{3})$) show that any pair of issues partitions the population into four consensus groups. Because the division of individuals into the groups (i.e., the 2-way consensus partition) is the same for any pair of issues, all pairs of issues exhibit the same amount of alignment. In other words, any alignment measure based on the resulting partition will yield the same score for all pairs of issues. However, when considering all three issues simultaneously, we find that the overall alignment of the system cannot be captured solely by examining the scores obtained at the pairwise level. Specifically, in this case, the consensus partition collectively induced by $T_{1}$, $T_{2}$, and $T_{3}$ does not preserve the group cleavages obtained from the $2$-combinations. This is because each pair of topics initially creates a larger consensus group of four individuals (taking the purple position for both issues), but the introduction of the third topic always ends up dividing this group into two smaller groups. Since the third topic is always cross-cutting one of the cleavages that exist at the pairwise level, the number of consensus groups increases, showing higher diversity of stances and lower 3-way alignment compared to what is observed at the pairwise level.
Second, relying solely on the assessment of pairwise alignment may hinder our ability to differentiate between different levels of higher-order alignment arising from the same pairwise relationships. We demonstrate this with Figure 2b, which shows the same $T_{1}$ and $T_{2}$ as in Panel (a), but changes $T_{3}$ to a different $T^{\prime}_{3}$. The systems in Panel (a) and in Panel (b) look identical at the pairwise level, as the consensus partitions $C(T_{1},T_{3})$ and $C(T_{1},T^{\prime}_{3})$, and $C(T_{2},T_{3})$ and $C(T_{2},T^{\prime}_{3})$ are equivalent. However, considering all three issues simultaneously yields different consensus partitions $C(T_{1},T_{2},T_{3})$ and $C(T_{1},T_{2},T^{\prime}_{3})$. Therefore, the systems in Panel (a) and in Panel (b) exhibit different levels of 3-way alignment, with the former being less aligned (i.e., having a greater number of consensus groups).
The possibility of observing different higher-order constraints arising from the same set of pairwise alignments underscores the necessity of extending beyond pairwise analysis, as relying solely on comparisons between pairs of topics poses the risk of inaccurately portraying the higher-order alignment of a system. By studying consensus groups defined by individuals’ opinions on multiple topics, we are able to define a measure of multiway alignment that captures the overall consistency of beliefs across an entire spectrum of political topics.
3 Measuring Multiway Alignment
In this section, we formally develop our multiway alignment measure. In Section 3.1, we formalize the idea of consensus partitions and its significance in investigating the extent to which opinion groups across multiple topics are either aligned or cross-cutting. In Section 3.2, we produce a list of desired properties for a multiway alignment measure and present our formulation for it.
3.1 Consensus Partitions
Formally, we investigate a set of individuals $I$, and how they are divided to opinion groups $g\subseteq I$ for a given topic. We assume that each individual holds exactly one opinion in a single topic, such that the opinion groups are mutually exclusive. That is, each topic $T$ can be represented as a partition of the set of all individuals $I$, which is formally defined as
$T\subseteq\{g\quad|\quad g\subseteq I\},\quad\text{ such that }\quad\emptyset\notin T,\quad g\cap g^{\prime}=\emptyset\quad\forall g\neq g^{\prime}\in T,\quad\text{ and }\quad\bigcup_{g\in T}g=I\,.$
To quantify alignment in the sense of Figure 1, we identify clusters of individuals that consistently belong to the same opinion group, i.e., are like-minded, across the given $k$ topics $T_{1},\ldots,T_{k}$. We refer to these clusters as consensus groups defined by the set $T_{1},\ldots,T_{k}$. The set of consensus groups is a consensus partition as each individual again belongs to exactly one consensus group. Formally, a consensus partition $C(T_{1},\ldots,T_{k})$ is defined as the set of non-empty intersections of the partitions defined by each topic:
$C(T_{1},\ldots,T_{k})=\left\{c\subseteq I\quad|\quad\exists(g_{1},\ldots,g_{k}),\quad g_{i}\in T_{i}\quad\text{s.t.}\quad\bigcap_{i=1}^{k}g_{i}=c,\quad c\neq\emptyset\right\},$ (1)
where $T_{i}$ represents the partition of individuals defining the opinion groups for the $i$-th topic.
Besides having a clear interpretation in terms of societal dynamics, the consensus partition also effectively represents the joint distribution of opinions across the topics under consideration. Consequently, the entropy $H(T_{1},\ldots,T_{k})=H(C(T_{1},\ldots,T_{k}))$ represents the amount of information needed to describe the consensus partition, which reflects the complexity of the identity space. When individuals mostly agree with their ingroup on various topics or divides, the consensus partition has low entropy: even when the set of topics is large, the amount of information required to describe the consensus partition is minimal, as there are only few entrenched consensus groups. Conversely, a consensus partition with high entropy describes a complex identity space with cross-cutting cleavages. Because of the multitude of diverse consensus groups, the identity space is more heterogeneous, and the consensus partition requires a longer description to capture all the identities.
The emergence of belief systems, then, can be characterized by the amount of information that each individual topic contains about the consensus partition obtained from the other topics in analysis. Specifically, in a highly interdependent system, the uncertainty in the consensus partition greatly decreases when we acquire knowledge about the opinion groups on one topic. Conversely, in the absence of constraint, knowing the opinion groups on one topic provides little explanation of the configuration of the consensus partition.
3.2 Multiway Alignment
We start building our multiway aligned measure by listing desirable properties, and use this list to construct the measure:
  1. 1.
    The measure should be able to describe higher-order phenomena while still having an intuitive interpretation and correspondence with the relevant theoretical concepts of consensus and opinion groups we developed above.
  2. 2.
    When the multiway alignment is computed for $n=2$, the multiway measure should reduce to an existing pairwise measure that could be used to quantify alignment. For example, our multiway alignment measure should reduce to the pairwise alignment measure used in Chen et al. (2021) if we define it in terms of normalized mutual information (NMI). This ensures consistency and comparability with existing methods.
  3. 3.
    The measure should be clearly normalized and not depend on factors such as the number of individuals or imbalance of the opinions. We would like the value 1 to indicate complete alignment and 0 to indicate no alignment. This ensures that the results are easily interpretable and comparable across different contexts.
  4. 4.
    The measure should be designed to avoid overfitting. It should clearly differentiate between true alignment and alignment that can be attributed to random chance, possibly by allowing for statistical significance test or a baseline comparison.
Property 1
Given the interpretation of the consensus partition presented in Section 3.1, we establish a quantitative measure of multiway alignment across a set of $k$ topics based on how similar the opinion partition of each topic is to the consensus partition of the $k-1$ remaining topics. In particular, for a given list of topic $T_{1},\ldots,T_{k}$, we can quantify their multiway alignment by averaging a chosen similarity measure S between each topic and the consensus partition defined by the remaining topics:
$A_{\text{S}}=\frac{1}{k}\sum_{i=1}^{k}\text{S}(T_{i},C(T_{1},\ldots,T_{i-1},T_{i+1},\ldots,T_{k})).$ (2)
If the similarity is based on mutual information, i.e., $S=MI$, the measure describes the average amount of information knowing someone’s opinion in one topic gives about their position in all of the other topics. As shown in D, $A_{\text{MI}}$ can be interpreted as the total correlation, adjusted by subtracting contributions from lower-order alignments.
Property 2
As long as the chosen S is symmetric, $A_{S}$ reduces to the original pairwise measure $S$ for $k=2$, since the consensus partition $C(T_{j})$ is defined by topic $T_{j}$ alone and therefore matches opinions in $T_{j}$ itself. By the same reasoning above, the formula in Equation 2 simplifies to exactly the measure used in Chen et al. (2021) for $k=2$ when S = NMI. Yet, the formulation in Equation 2 allows us to quantify also alignment across multiple topics, as the mutual information is now computed between each topic and the consensus partition defined by the remaining topics in analysis.
Property 3
By definition, normalized mutual information (Vinh et al., 2009) takes on values between 00 and $1$, where 00 indicates independence and $1$ signifies complete agreement between the two labeling assignments that are compared. The $A_{NMI}$ score naturally inherits those properties, and therefore allows for similar interpretation. NMI, however, does not correct for agreement solely due to chance and thus may yield inflated scores, particularly when comparing groups with significant size imbalances (Newman et al., 2020; Salloum et al., 2022). This limitation also applies to the multiway alignment score $A_{NMI}$. We mitigate this issue by using the adjusted mutual information (AMI) (Vinh et al., 2009) instead, and all our numerical results will be presented in terms of $A_{AMI}$.
Property 4
We validate the statistical significance of our multiway alignment findings by introducing a null model for comparison (F). The null model provides the expected outcome $\langle A_{\text{null}}\rangle$ when the multiway alignment is a consequence of random chance and allows us to establish confidence intervals that account for fluctuations in the value of $A$ due to the finite number of individuals in the system.
A multiway alignment score defined as in Equation 2 with $S=AMI$ satisfies the four properties we stated above. Additionally, we define the maximal alignment curve (C) to describe the evolution of alignment as an increasing number of topics are considered. The maximal alignment curve collects the maximum alignment scores reached by the system at any possible combinations of $k$ topics. By considering the sets of topics exhibiting the highest alignment, the shape of the maximal alignment curve summarizes the overall alignment of the system, and the area under this curve can be used as a global summary statistic, describing the higher-order alignment of the system.
4 Applications
Figure 3: The shaded areas in Panel (a) show the range of values obtained for the multiway alignment across 43 different systems, while the lines represent the maximal alignment curves for selected periods and systems. Panel (b) and (c) show the spectrum of multiway alignment in Finnish Parliament over the parliamentary years 2018/2019 and 2020/2021 respectively. Panel (d) shows the spectrum of the multiway alignment obtained from the responses to ANES surveys in 2020. The spectrum in Panel (e) is obtained from the retweet networks before Finnish elections 2023. Panel (f) considers the votes in the U.S. House in 2014.
In this section, we illustrate higher-order alignment patterns revealed by our multiway alignment measure in case studies that cross national, institutional – public opinion surveys, political chambers, and social media platforms – and temporal contexts. Specifically, we analyzed Finnish Parliament votes from 1996–2021, ANES surveys collected every four years from 2004–2020, pre-election discussions on Twitter in 2019 and 2023, and U.S. House activities from 2004–2013. Figure 3 summarizes all results from these systems and years. The first panel, Figure 3a, shows with the shaded areas the range of multiway alignment values obtained from the 43 systems we analyzed, and examples of the maximal alignment curve. The remaining panels show the full spectrum of multiway alignment corresponding to each maximal alignment curve in Figure 3a. The spectrum includes all scores obtained when considering sets of topics of size two and above, and highlights values that are significant at $\alpha=5\%$.
Despite the distinct nature of each system, the multiway alignment spectrum observed from our datasets generally tends towards a funnel shape (Figure 3b-e-f and corresponding maximal alignment curves in Figure 3a for Finnish Parliament 2018–2019, Finnish Twitter 2023 and U.S. House 2014). Yet, the shape of the spectrum can vary substantially: the difference in Figure 3b-c and in the maximal alignment curves for Finnish Parliament 2018–2019 and 2020–2021 (Figure 3a) reflect a well-recognized impact of the pandemic, that we further discuss in Section 4.1.
Figures 3b-c (Finnish Parliament) and Figure 3f (U.S. House) demonstrate how parliamentary and governmental chambers often exhibit high multiway alignment due the collaborative nature of those legislative assemblies, where alignment reflects teamsmanship and elite polarization. Conversely, the level of higher-order alignment among the general public is lower, and further decreases when the sample of the population is random. This is evident from Figure 3a, where the values obtained when measuring alignment from ANES surveys are clearly the lowest among the four systems. Still, as discussed in Section 4.2 and Section 4.3, we observe an increasing trend in alignment also in the general population.
4.1 Multiway Alignment in Finnish Parliament Before and During COVID Pandemic
Figure 4: Multiway alignment in Finnish Parliament over time, detailing the patterns before and during the COVID pandemic (2019–2020 and 2020–2021).
The analysis of multiway alignment within the Finnish Parliament offers insights into the shifting political dynamics before and during the COVID-19 pandemic, and reveals a known phenomenon that purely pairwise analysis would miss. The line chart in Figure 4 shows the evolution over time of the overall alignment, summarized through the area under the maximal alignment curve. The two points highlighted in the line chart show how the emergence of the pandemic and the subsequent declaration of a state of emergency led to a sharp decrease in alignment in 2020–2021.
A comparison of the multiway alignment spectrum for the 2019–2020 and 2020–2021 parliamentary periods is shown in the main panel in Figure 4. Before the onset of the pandemic, the spectrum of the multiway alignment within the Finnish Parliament displayed a funnel shape that we commonly observe in legislative systems, where the alignment across different topics reflects the dynamics of policymaking and political discourse. The funnel shape of multiway alignment indicates a gradual decrease in alignment as the number of topics in the analysis increases. Nevertheless, the values can remain at values above zero also at higher orders (e.g., $k=9$ and $k=10$ in 2019–2020), demonstrating alignment that could originate from adherence to party lines, whereby members from the same party (or coalition) vote in a similar way on a wide range of issues. [^1] The same effect is also visible in Figure 3f, which shows one instance of the alignment spectrum for U.S. House (Appendix G). Being a two-party system, the U.S. witnesses a much more pronounced effect of elite polarization on the spectrum of multiway alignment, and the multiway alignment score remains at moderate values of around 0.3 even when considering all the topics discussed over the year ($k=18$).
Notably, during the COVID-19 pandemic, especially early on, Finnish parliamentary politics exhibited unconventionally high levels of cooperation between the government and opposition coalitions (Lehtonen and Ylä-Anttila, 2024). This was evident in how both sides repeatedly appealed for broad cooperation, reciprocally praise each other for implemented COVID responses, and even in how the opposition withdrew their interpellation on the government’s immigration policy (Eduskunta.fi, 2020). Despite this drastic shift from conventional politics, the overall patterns gleaned from a pairwise alignment analysis exhibit no visible difference across the two periods. The only hint of a reorientation of priorities and focal points within parliamentary deliberations is a shift in the most aligned pair of topics, from defense and health in 2019–-2020 to transportation and health in 2020-–2021.
The changes brought by the pandemic become glaringly evident only when analyzing the patterns of alignment at higher orders: the sharp drop in multiway alignment following $k=7$ and the manifest shift in the distribution of the scores already at $k=3$ usher in a period of exceptional policy discussions. At $k=3$, the spectrum for 2020–2021 shows two clear clusters of topics, one retaining alignment levels comparable to the previous period, and the other exhibiting markedly low alignment. The same clustering effect becomes increasingly distinct at higher orders ($k=4,5,6$). Overall, the spectrum of alignment in 2020–2021 clearly shows six topics – transportation, culture, defense, economics, foreign affairs, and health – following similar alignment patterns as in 2019-2020. The remaining five topics – climate, education, immigration, social issues, and unemployment – become much less aligned in 2020–2021, bringing the multiway alignment score at $k=10$ and $k=11$ towards zero.
Looking more deeply into these patterns, we see that they reflect the collaborative dynamics explicitly adopted to face the crisis. On one hand, the sudden prominence of topics related to transportation and their alignment with health and economics suggests a focused effort within the Finnish Parliament to navigate the circumstances imposed by the pandemic through initiatives related to free movement restrictions, economic stimulus packages, quarantine and isolation protocols, and wide vaccination efforts (YLE, 2020a, b). On the other hand, topics such as climate and immigration could have been perceived as less immediately relevant or lower in priority compared to the urgent challenges posed by the crisis. Overall, our findings quantitatively support the observation that, throughout the pandemic, parliamentary discussions diverged from traditional political dynamics, and that the opposition adopted a collaborative approach, refraining from strong criticism of COVID policies proposed by the government (Niemikari and Raunio, 2022).
4.2 Evolution of Alignment in Finnish Twitter Before Elections
Figure 5: The comparison of multiway alignment in Finnish Twittersphere before 2019 and before 2023 elections shows an overall increase in alignment, as well as changing patterns in the alignment induced by issues and parties discussed before elections.
Next, we explore multiway alignment patterns in Twitter discussions in the months leading up to the 2019 and the 2023 Finnish political elections. We show these results in Figure 5. The spectrum of the multi-alignment across selected salient issues (left panel in Figure 5) and party preferences for the major five parties (right panel in Figure 5) unveils evident trends and shifts in inter-topic and inter-party relationships.
Both panels in Figure 5 confirm the previously observed significant rise in pairwise alignment in the Finnish online discourse (Salloum et al., 2024). However, they also show a concerning increase in higher-order alignment across all $k$, underscoring a heightened interdependency among political elements discussed online. Exemplification of this are the instances of $k=3$ and $k=4$ in Figure 5 (left panel), where the least aligned $k$-tuple of issues in 2023 surpasses the alignment of the most aligned $k$-tuple of issues in 2019, indicating a substantial increase in alignment as measured through political discussions on Twitter.
While an increase in higher-order alignment is also evident when analyzing party positions, the magnitude of this rise is notably smaller compared to the trends observed among issues. The emergence of alignment within discussions around political parties, particularly in a multi-party system where coalition governments are commonplace, is somewhat anticipated. However, the substantial upsurge in higher-order alignment across issues signifies a shift towards a more partisan-centric political landscape. Moreover, the difference in the evolution of alignment over time between issues and party preferences suggests a transformation in the nature of political discourse. The increasing alignment among issues indicates a trend where discussions are increasingly influenced by partisan ideologies rather than a mere representation of varied political party preferences. This shift might imply a growing polarization of political viewpoints across a diverse range of issues, emphasizing the intertwining of diverse concerns into a more ideologically driven political sphere.
Figure 5 also reveals shifts in issue relevance over time. In 2019, the most aligned pair of issues comprised economic policy and immigration, whereas in 2023, immigration and social security emerged as the most aligned pair, with economic policy replacing climate change in the most aligned 3-tuple. The analysis of discussions around parties in 2019 hint at a constraint between the opinion of Twitter users on SDP, the election winner, and the National Coalition Party. In contrast, the online discourse surrounding 2023 elections showcases a correlation between the perspectives on the Finns Party and the Green Party, which occupy opposing roles in the current government. Beyond government-opposition dynamics, this shift may also reflect a broader transition from the traditional left-right political identities that were dominant in 2019 to a nationalist-internationalist divide by 2023.
4.3 Multiway Alignment and Party Preferences: Higher-Order Partisan Sorting of the U.S. Public
Figure 6: A comparison of multiway alignment between ANES Time Series data 2004 and 2020.
Finally, we study multiway alignment (both issue alignment and partisan alignment) in U.S. public opinion using the ANES Time Series data from 2004–2020. First, we show in Figure 6 the full spectrum of the multiway alignment computed from the 2004 and 2020 data, which reveals a shifting salience of relationships over this temporal span among eight topics (seven issues and partisan identification). In particular, Figure 6 shows a transition from the predominant alignment of stances concerning abortion and LGBT rights in 2004 to a more pronounced alignment between gun access and immigration in 2020. Notably, the alignment score pertaining to the latter pair surpasses the three highest scores observed in 2004, signifying the existence of a stronger 2-way constraint in 2020. Beyond this pairwise analysis, Figure 6 also shows an increase over time in the overall multiway alignment across all eight topics, indicating intensified interdependencies among the analyzed stances and portraying a more polarized landscape of the contemporary American discourse.
The analysis of 3-tuples and higher illustrates the growing importance of partisanship in U.S. politics. Whereas in 2004, positions toward immigration played a central role within triadic-and-higher relationships, in 2020, party preferences became one of the most prominent topics, marking a shift towards sociopolitical affiliations as key alignment determinants. In this respect, our measure of multiway alignment allows us to investigate research questions that study partisanship across a multitude of issues. Specifically, we are now able to quantify the strength of the relationship between party preference and a bundle of multiple topics.
We show this in the heatmap in Figure 7, which specifically illustrates the influence of including party preference into any 3-tuple formation as the percentage change in multiway alignment score. Notably, starting from 2012, the 4-way alignment of any four topics that include individual party preferences consistently surpasses the corresponding alignment scores obtained from the three topics other than party preference. This phenomenon is not restricted to 3-tuples: the relative change of the multiway alignment score when incorporating individual party preference within $k$-tuples ($k\in\{2,\dots,6\}$) over a more complete range of years shows a visible trend (Figure 7), demonstrating an increasing alignment over time between party preferences and the remaining stances, across all values of $k$. The observed development of multiway alignment confirms the growing entwinement between political affiliations and multiple issues, suggesting that stances on various issues are increasingly catalyzed by political affiliations. These results not only demonstrate an increasing trend in partisan alignment across the general public of the U.S., but also identify the period between 2008 and 2012 as a pivotal point where the multi-dimensional constraint between party affiliation and the set of issues considered in this study becomes clearly significant.
Figure 7: The heatmap shows the percentage change of multiway alignment when adding party preference to any $3$-tuple of topics selected from ANES Time Series Study (American National Election Studies, 2021). The color gradient shows the increasing significance of party preference over time in $4$-way alignment. The timeseries chart in the bottom right shows how the same trend holds for tuples of any size.
5 Discussion
We build upon existing approaches for the study of pairwise alignment and provide a quantitative method to analyze alignment across multiple topics. Our proposed approach retains similar interpretability as its corresponding pairwise version, enabling the investigation of issue alignment and partisan alignment across various systems. We demonstrated its utility in contexts where opinion groups are straightforwardly defined by pre-existing categories, such as surveys. We also illustrated its applicability in instances where individual group membership is not explicitly available, such as when looking at opinion groups observed in online social networks. Additionally, when applied to Parliamentary data, our multiway alignment measure identifies subsets of issues that exhibit high alignment based on co-voting patterns and irrespective of party affiliation.
Finally, we note that the sample size required for the analysis of many topics can be a limiting factor, as multiway analysis becomes more meaningful when the size of the sample of individuals substantially exceeds the product (across topics) of the number of opinions that can be manifested at each topic.
Our work highlights how multiway alignment provides valuable insights into the dynamics of socio-political systems, revealing the presence of higher-order constraints across various contexts. While our case studies represent only a subset of the potential applications of our method, they demonstrate its accuracy and robustness. Looking ahead, our work sets the stage for future research on societal polarization and the emergence of partisan animosity. Additionally, our research equips researchers with computational tools, made available through the open-source Python package multiway_alignment, to quantify and visualize multiway alignment as we have in this paper.
Acknowledgement
We thank Boyoon Lee and Kevin Reuning for helpful feedback.
Funding Statement
This research was supported by grants from the Research Council of Finland (349366; 352561; 357743)
Competing Interests
None.
Data Availability Statement
Replication code for this article is available at https://github.com/letiziaia/multiway-alignment. All data used in this study is available publicly. For details, see Appendix A and Appendix B.
Appendix
Appendix A Data
We operationalize the suggested multiway alignment framework across diverse datasets, including survey data (described in Appendix A.1), retweet networks (Appendix A.2), and Parliamentary roll-call votes (Appendix A.3).
A.1 Survey Data
The survey data are sourced from the ANES 2020 Time Series Study, encompassing interview responses from eligible U.S. voters across various years. The interview questions span a spectrum of issues, including topics such as public health, faith in experts and science, group empathy, and societal concerns. From the available range of topics, we selected questions about immigration (VCF9223), abortion (VCF0838), gun access (VCF9238), unemployment (VCF9229), party preference (VCF9205), death penalty (VCF9236), LGBT rights (VCF0876), government spending (VCF0839). This selection facilitates an examination of the evolution of multi-issue alignment across the years 2004 to 2020.
A.2 Retweet Networks
In line with the methodology employed in (Chen et al., 2021), we apply our algorithm to Twitter data related to Finnish political elections in 2019. Additionally, we extend our analysis to include data from the Finnish political elections in 2023 (Salloum et al., 2024).
For the 2019 dataset, tweets and retweets were gathered during the same period (March to July 2019), using a predefined set of Finnish hashtags as in (Chen et al., 2021). The collected data were then categorized into 11 distinct labels based on their text content. These labels encompass conversation streams related to Finnish political parties (SDP, National Coalition, Finns Party, Green Party, Left Party, Center Party) and salient topics (economic policy, social security, immigration, climate, education). Given that a single tweet can relate to multiple categories simultaneously, this approach results in a multilabel classification. Subsequently, for each label, we constructed the corresponding retweet network as detailed in Appendix B.2.
The dataset related to the 2023 Finnish elections comprises tweets and retweets published from September to April 2023, and was collected by using a similar set of Finnish hashtags as for 2019 (Chen et al., 2021). Similar to the 2019 data, we categorized tweets and retweets into the same 11 labels and subsequently build the retweet networks by following the same approach as for 2019.
A.3 Roll-Call Data
The multiway approach described in the paper has been applied to voting data in Finnish Parliament and U.S. House. Data sources and data description are provided below.
A.3.1 Finnish Parliament
The dataset of Parliamentary activities has been collected via the REST API offered by the Finnish government (edu, 2022). The data is available both in Finnish and in Swedish, but policy topics are not explicitly given. As the analysis has been conducted on the Finnish version of the data, the Finnish titles of the bills have been used to derive the policy labels. In order to do so, Finnish words have been stemmed and stop-words have been removed. Then, we obtain a feature matrix with latent semantic analysis (LSA), by applying truncated singular value decomposition (SVD) to the term frequency-inverse document frequency (tf-idf) matrix. Finally, major topics are found by K-Means clustering, where the number of clusters is chosen by minimizing Davies-Bouldin score (Davies and Bouldin, 1979). The Finnish version of the data from edu (2022) indicates each Parliament member’s vote by utilizing one among jaa (yes), ei (no), poissa (absent), tyhjä (empty). For the further analysis, the voting behaviour is simplified to yes, no, and other, where yes and no correspond exactly to the original categories, whereas the category other covers both poissa and tyhjä. Among the 33 different parties that appear in the Finnish Parliament between 1996 and 2021 (Figure 8), the 8 biggest parties were used for the analysis as they appear in all years and account for the vast majority of the Parliament. Specifically, the 8 parties included in the analysis are Centre Party (kesk), National Coalition Party (kok), Finns Party (ps), Swedish People’s Party of Finland (r), Social Democratic Party (sd), Left Alliance (vas), Green League (vihr), Christian Democrats (kd). For the following analysis, each Parliamentary period is divided into 4 approximately equal-sized sub-periods.
Figure 8: Number of Parliament Members (MPs) by party in the Finnish Parliament.
A.3.2 U.S. House
A complete dataset covering all congresses from 1 to 117 and both chambers (House and Senate) is available from Voteview (usd, 2021). As for policy topics, the ones provided by Comparative Agendas (Adler and Wilkerson, 2017) for congresses between 97 and 113 are linked to the corresponding votes. Since the Independent party (party code 328) has very limited number of members over the Congresses, it has been omitted from the analysis. As a result, the findings describe the dynamics of members of Republican Party and Democratic Party only. Furthermore, each Congress is split into 2 approximately equal-sized sub-periods, roughly corresponding to calendar years.
Appendix B Opinion Partitions from Clustering or Community Detection
Opinion groups have been extensively studied using clustering and community detection techniques (Conover et al., 2011; Chen et al., 2021; Brabec, 2020). These are powerful techniques for uncovering patterns in data, identifying distinct subgroups based on similarities in their actions or beliefs, and shedding light on the dynamics of polarization or on the diffusion of ideas within a society or network. Taking a similar approach, the idea of clusters or communities can be harnessed to investigate opinion groups across multiple issues simultaneously.
B.1 Parliamentary Roll-Calls
In a roll-call vote, each member of the Parliament can vote either yes or no or can abstain. However, for each policy area, there are typically multiple roll-call votes. Therefore, for each policy area, we can infer the opinion groups from the similarity in the vector of votes each individual MP has casted on all the roll-calls regarding a specific policy domain.
For each issue or policy domain, each of the Parliament members (MPs) is represented with the vector of votes they have casted on all the roll-calls regarding the specific issue. Specifically, for each issue, each MP is associated to a vector $u_{i}=(u_{i1},u_{i2},...)$, where the element $u_{ij}$ is the vote casted by the MP $i$ on the $j$-th roll-call regarding that issue, and can take values $-1$ (no), $1$ (yes), or 00 (other, as described in Appendix A.3). This representation allows us to detect MPs with highly co-voting behavior from the cosine similarity $u_{m}.u_{n}$ between their vectors. In fact, Parliament members who cast the same votes have small cosine distance and high cosine similarity. As a result, for each issue, we can build the similarity network of the members of Parliament, where the interaction between each pair of MPs is modeled by a link weighted proportional to the similarity of their voting behavior. More precisely, given one policy domain or issue, the network of Parliament members is defined as follows:
  • each MP is mapped to a node in the network;
  • each pair of MPs is connected by a link that is weighted by the cosine similarity between the vectors of votes casted by the two MPs, where each vote is either $1$ (yes), $-1$ (no), or 00 (has not voted).
By definition, the resulting network is an undirected weighted all-to-all network that carries information about co-voting behavior of the MPs on one specific issue. The remaining layers of the network are defined similarly, each time by considering the votes associated with the topical issue of interest.
Finally, we repeatedly apply clustering to each issue network independently, to find communities of strongly collaborating MPs at each layer. Specifically, the opinion partitions for each issue for Parliamentary roll-calls data are discovered via layer-level clustering performed with Density-Based Spatial Clustering of Applications with Noise (DBSCAN) (Ester et al., 1996), that detects areas of high density, separated by areas of lower density. Therefore, the algorithm neither requires to know the number of clusters beforehand nor assumes convexity of the clusters. The clusters are expanded from a number of core samples by recursively including their neighbours. On the other hand, samples that are distant from any core points are considered to be outliers. In practice, this means that, for each policy domain or topic $T_{i}$, the algorithm is able to find the optimal number of clusters and a number of outliers, depending on the co-voting behaviour on the particular issue.
The hyperparameters of DBSCAN are optimized based on the Silhouette score (Rousseeuw, 1987), which compares the mean intra-cluster distance and the mean nearest-cluster distance to assess how well the clusters are separated. By applying DBSCAN to the pre-computed cosine distance matrix of voting behavior, we identify groups of highly collaborating MPs over each issue. Therefore, the labelings resulting from DBSCAN define the opinion partitions on which we apply the proposed multiway alignment measure.
B.2 Retweet Networks
In each of the retweet networks, nodes represent users who shared a tweet (original or retweet) containing at least one relevant hashtag. Nodes are connected by an undirected edge if one user has retweeted the other on a given topic. Following the approach established in (Chen et al., 2021) and (Garimella et al., 2018), we employed the METIS algorithm (Karypis and Kumar, 1998) to separately partition each networks into two similarly-sized groups with minimal connections between them. To maintain a realistic representation of group distribution in the network, a maximum imbalance constraint of 3:7 was applied, preventing the algorithm from producing a trivial partition. Partitions for each of the retweet networks are available in Zenodo (Iannucci et al., 2024).
Appendix C Maximal Alignment Curve
The score defined in Equation 2 can be computed for each $k$-combination of issues. For every value of $k$, our primary interest lies in the $k$-combination that maximizes $A$. To this scope, the maximal alignment curve summarizes the behavior of the system by considering the sets of issues exhibiting the highest alignment. In fact, the maximal alignment curve is a collection of the highest multiway alignment scores at each $k$.
As the maximal alignment curve is upper-bounded by $1$, the area under it describes the alignment of the entire system across all the issues. In fact, if the multiway alignment score remains close to $1$ across all $k$, it signifies that all issues exhibit extraordinarily high alignment. Conversely, the maximal alignment curve rapidly decreases if the consensus groups deteriorate, degenerating towards singletons as more issues are considered.
1
2function getAlignmentCurve $(M)$;
Input : A 2-dimensional matrix $M$ of size $n\times m$, where $n$ is the number of individuals, $m$ is the number of issues, and $M_{i,j}$ is the label of the opinion group to which person $i$ belongs with respect to issue $j$
3
Output :  getAlignmentCurve$(M)$, mapping each maximally aligned subset of issues of size $2,\dots m$ to its $A_{S_{k}}$ score
4
5result = Map();
6for $k\leftarrow 2$ to $m$ do
7
8      allCombinationsSizeK = getCombinations(issues, k);
9      bestScore = 0;
10      for $c\leftarrow 1$ to ($m\choose k$ do
11
12            issueSubset = allCombinationsSizeK[c];
13            topicsOpinions = M[$1:$n][issueSubset];
14            consensusPartition = getConsensusPartion(topicsOpinions);
15            score = 00;
16            for $j\leftarrow 1$ to k do
17
18                  score +=+=+ = S(consensusPartition, topicsOpinions[$1:$n][j]);
19             end for
20
21            score /=/=/ = k;
22            if score $>$ bestScore then
23
24                  bestScore = score;
25                  bestKCombination = policies;
26             end if
27
28       end for
29
30      result[bestKCombination] = bestScore;
31 end for
32
return result;
Algorithm 1 Maximal alignment curve
Appendix D Interpretation of $A_{MI}$ Measure
Given the set of $n$ topics $\mathcal{N}=\{T_{1},\ldots,T_{n}\}$, we show that, when the similarity measure $S$ is the mutual information (MI), the definition of multiway alignment given in Equation 2 is equivalent to
$A^{(n)}=\left(\sum_{i=1}^{n}H(T_{i})\right)-H(T_{1},\ldots,T_{n})-\frac{1}{n}\left(\sum_{s\in{\mathcal{N}\choose n-1}}\sum_{k=2}^{n-1}\bar{A_{s}}^{(k)}\right)$ (3)
where the terms $\bar{A_{s}}$ are averages of lower-order multiway alignment over subsets of $\mathcal{N}$.
By our construction in Section 3.2,
$A^{(n)}=\frac{1}{n}\sum_{i=1}^{n}\text{MI}(C(T_{1},\ldots,T_{i-1},T_{i+1},\ldots,T_{n}),T_{i}),$
where MI is the pairwise mutual information, $C(T_{1},\ldots,T_{i-1},T_{i+1},\ldots,T_{n})$ is the consensus partition of individuals defined by the topics $T_{1},\ldots,T_{n}$ excluding $T_{i}$, and each $T_{i}$ defines the opinion groups for the $i$-th individual topic. Therefore, for $n=2$, $A^{(2)}=MI(T_{1},T_{2})=\sum_{i=1}^{2}H(T_{i})-H(T_{1},T_{2})$. Similarly, for $n=3$,
$A^{(3)}=\left(\sum_{i=1}^{3}H(T_{i})\right)-H(T_{1},T_{2},T_{3})-\frac{1}{3}\left({A^{(2)}}_{23}+{A^{(2)}}_{13}+{A^{(2)}}_{12}\right),$
as $\bar{A_{s}}^{(n-1)}={A_{s}}^{(n-1)}$ for each $s\in{\mathcal{N}\choose n-1}$.
Let us now assume that the statement in 3 is true for some arbitrary $n-1\geq 2$, that is,
$A^{(n-1)}=\left(\sum_{i=1}^{n-1}H(T_{i})\right)-H(T_{1},\ldots,T_{n-1})-\frac{1}{n-1}\left(\sum_{s\in{\mathcal{N}-1\choose n-2}}\sum_{k=2}^{n-2}\bar{A_{s}}^{(k)}\right)$
and therefore
$H(T_{1},\ldots,T_{n-1})=\left(\sum_{i=1}^{n-1}H(T_{i})\right)-\frac{1}{n-1}\left(\sum_{s\in{\mathcal{N}-1\choose n-2}}\sum_{k=2}^{n-2}\bar{A_{s}}^{(k)}\right)-A^{(n-1)}.$
Then,
$A^{(n)}=\frac{1}{n}\sum_{i=1}^{n}\text{MI}(C(T_{1},\ldots,T_{i-1},T_{i+1},\ldots,T_{n}),T_{i})$
$=\frac{1}{n}\sum_{i=1}^{n}H(T_{i})+\frac{1}{n}\sum_{i=1}^{n}H(T_{1},\ldots,T_{i-1},T_{i+1},\ldots,T_{n})-H(T_{1},\ldots,T_{n})$
$=\frac{1}{n}\sum_{i=1}^{n}H(T_{i})+\frac{1}{n}\sum_{i=1}^{n}\left(\sum_{j=1,j\neq i}^{n}H(T_{j})-\frac{1}{n-1}\left(\sum_{t\in{\mathcal{N}-1\choose n-2}}\sum_{k=2}^{n-2}\bar{A_{s}}^{(k)}\right)-A^{(n-1)}\right)-H(T_{1},\ldots,T_{n})$
$=\frac{1}{n}\sum_{i=1}^{n}H(T_{i})+\frac{1}{n}\sum_{s\in{\mathcal{N}\choose n-1}}\left(\sum_{j\in s}H(T_{j})-\frac{1}{n-1}\left(\sum_{t\in{s\choose n-2}}\sum_{k=2}^{n-2}\bar{A_{t}}^{(k)}\right)-{A_{s}}^{(n-1)}\right)-H(T_{1},\dots,T_{n})$
$=\frac{1}{n}\left(\sum_{i=1}^{n}H(T_{i})+\sum_{i=1}^{n}\sum_{j=1,j\neq i}^{n}H(T_{j})\right)-\frac{1}{n}\sum_{s\in{\mathcal{N}\choose n-1}}\left(\sum_{k=2}^{n-2}\bar{A_{s}}^{(k)}+{A_{s}}^{(n-1)}\right)-H(T_{1},\dots,T_{n})$
$=\sum_{i=1}^{n}H(T_{i})-\frac{1}{n}\left(\sum_{s\in{\mathcal{N}\choose n-1}}\sum_{k=2}^{n-1}\bar{A_{s}}^{(k)}\right)-H(T_{1},\ldots,T_{n}),$
which also shows that each ${A_{s}}^{(k)}$ term contributes to the score by a fraction $\frac{1}{n\cdot(n-1)\cdot\ldots\cdot(k+1)}$ as long as $3\leq k+1\leq n$. For example, when $n=4$,
$A^{(4)}=\left(\sum_{i=1}^{4}H(T_{i})\right)-H(T_{1},T_{2},T_{3},T_{4})-\frac{1}{4}\left(A^{(3)}_{123}+A^{(3)}_{134}+A^{(3)}_{124}+A^{(3)}_{234}\right)-\frac{1}{4}\left(\bar{A}^{(2)}_{123}+\bar{A}^{(2)}_{124}+\bar{A}^{(2)}_{134}+\bar{A}^{(2)}_{234}\right)$
where
$\bar{A}^{(2)}_{123}=\frac{1}{3}\left(A_{23}+A_{13}+A_{12}\right)$
and each term of order $3$ contributes to the score by $\frac{1}{4}$. Therefore, the term
$\frac{1}{n}\left(\sum_{s\in{\mathcal{N}\choose n-1}}\sum_{k=2}^{n-1}\bar{A_{s}}^{(k)}\right)$
in Equation 3 can also be understood as
$\frac{1}{n}\sum_{s_{1}\in{\mathcal{N}\choose n-1}}\left({A_{s_{1}}}^{(n-1)}+\frac{1}{n-1}\left(\sum_{s_{2}\in{s_{1}\choose n-2}}\left({A_{s_{2}}}^{(n-2)}+\frac{1}{n-2}\left(\sum_{s_{3}\in{s_{2}\choose n-3}}\left({A_{s_{3}}}^{(n-3)}+\frac{1}{n-3}\left(\ldots\right)\right)\right)\right)\right)\right).$ (4)
This average (avg) can be expressed recursively as:
$\text{avg}_{n}=\frac{1}{n}\sum_{s_{1}\in{\mathcal{N}\choose n-1}}\left({A_{s_{1}}}^{(n-1)}+\text{avg}_{n-1}(s_{1})\right),$
where
$\text{avg}_{n-k}(s_{k})=\frac{1}{n-k}{A^{(n-k-1)}_{s_{k}}}$
when $n-k=3$ and $n-k-1=2$.
Appendix E Relation to Other Multivariate Measures
Co-information (Bell, 2003) quantifies the amount of information gained about one random variable by observing the other variables:
$I(X;Y;Z)=I(X;Y)+I(X;Z|Y)$
$=H(X)+H(Y)+H(Z)-H(X,Y)-H(X,Z)-H(Y,Z)+H(X,Y,Z)\,,$
where $I(X;Y)$ is the mutual information between $X$ and $Y$, while $H$ is the entropy. In other words, co-information (Bell, 2003) captures how much information is collectively shared among all variables in a system, by quantifying the extent to which the joint distribution of all variables contributes to our knowledge about individual variables. Even though co-information is a generalization of mutual information, its interpretation differs in that this measure can also be negative in case of $3$ or more variables. As a non-negative generalization of mutual information, total correlation (TC) (Watanabe, 1960) is defined as the difference between the sum of the individual entropies of the variables and their joint entropy:
$TC(X_{1},\dots,X_{n})=\sum_{i=1}^{n}H(X_{i})-H(X_{1},\dots,X_{n}).$
In practice, TC measures the degree of interdependence among a set of variables by comparing their total entropy with the entropy of the joint distribution. In equivalent terms, this corresponds to measuring the amount of information that each individual variable contributes to the collective information represented by the joint distribution. Therefore, TC would be 0 if and only if the variables are independent. On the other hand, TC is maximized when all variables are completely determined by one of them. Similarly, dual total correlation (DTC) (Sun, 1975) or binding information (Abdallah and Plumbley, 2012) is non-negative, but, instead of being bounded by the sum of the entropies, it is bounded below by 0 and above by the joint entropy:
$DTC(X_{1},\dots,X_{n})=H(X_{1},\dots,X_{n})-\sum_{i=1}^{n}H(X_{i}|X_{1},\dots,X_{i-1},X_{i+1},X_{n}).$
Building on these measures, O-information (Rosas et al., 2019) or enigmatic information (James et al., 2011) quantifies higher-order interdependencies by the difference between TC and DTC, which can therefore result in negative scores for synergistic systems:
$\Omega(X_{1},\dots,X_{n})=TC(X_{1},\dots,X_{n})-DTC(X_{1},\dots,X_{n}).$
While those measures offer valuable insights, a range between 0 (no constraint) and 1 (total dependence) would represent an intuitive scale for a measure of higher-order constraint, helping the interpretability and usability of the measure across different contexts. To this end, DTC can be normalized to be in the range $[0,1]$ by dividing by the joint entropy. Additionally, it might be desirable for a measure of higher-order dependence to reduce to a pairwise measure of dependency in case there are only two entities involved. Among the measures reviewed above, this is true for co-information, which naturally reduces to the mutual information in case it is applied to only two variables.
As for the measure defined in Equation 2, it involves the individual entropies as well as the entropy of the joint distribution. In this sense, our measure extends the total correlation by considering also the contribution of lower-order dependencies. Furthermore, our formulation assesses the degree of interdependence among the variables by estimating on average how much information each variable contributes to the collective understanding of the system. Finally, our measure returns 0 in case of no constraint and 1 in case of total dependence.
Appendix F Null Model
The null model is defined as follows: for each topical issue in the system under examination, the number and sizes of opinion groups remain fixed, but the individual memberships are randomly permuted. This random permutation effectively creates a system with properties similar to the one under analysis. By averaging the results over multiple iterations of the null model, we obtain the expected alignment scores $\langle A_{\text{null}}\rangle$ for each $k$, and the expected maximum alignment curve for the system. These expected values account for multiway alignment purely attributable to chance.
Consequently, evaluating the multiway alignment scores $A$ (Equation 2) obtained from the real-life system involves comparing them to the expected multiway alignment scores $\langle A_{\text{null}}\rangle$ derived from the null model. Specifically, if the multiway alignment scores obtained from the real-life system fall within the $(1-\alpha)\%$ confidence interval of the scores generated by the null model, we can conclude that the alignment observed in the system is not statistically significant at the significance level $\alpha$. Conversely, the greater the difference between the multiway alignment score from the real-life system and the $(1-\frac{\alpha}{2})$-th percentile of the scores obtained from the null model, the stronger the evidence supporting the presence of high alignment in the system.
Therefore, we can adjust the multiway alignment measure defined in Equation 2 to account for the effect attributed to chance. To achieve this, we follow a similar logic as in Salloum et al. (2022) and subtract the expected alignment scores obtained from the null model from the observed multiway alignment scores in the real-life system. This subtraction effectively removes the influence of chance alignment. Subsequently, to ensure that the adjusted measure of alignment falls within the $[0,1]$ range, we renormalize it by dividing by $1-\langle A_{\text{null}}\rangle$.
The net multiway alignment score
$A_{net}=\frac{A-\langle A_{\text{null}}\rangle}{1-\langle A_{\text{null}}\rangle}$ (5)
serves as a robust and interpretable metric for quantifying the alignment of opinion groups across multiple topics, enhancing our ability to discern statistically significant alignment patterns within complex systems. The net alignment score represents, for each $k$, the excess alignment that cannot be attributed to chance, thus eliminating the implicit dependency of the multiway alignment score on the number of topics $k$.
Our empirical results show that $\langle A_{\text{null}}\rangle$ can be far from 0 when utilizig normalized mutual information (NMI), whereas $\langle A_{\text{null}}\rangle$ is consistently close to 0 when using AMI.
Appendix G Multiway Alignment of Elites: U.S. House
Figure 9: The spectrum of alignment measured from U.S. House in 2014 shows clusters of $k$-tuples that all reach high level of multiway alignment, as well as a value of higher-order alignment at $k=18$ still very far from 0.
Both the House and Senate exhibit clear patterns of alignment that could originate from party lines within these legislative chambers. The alignment spectrum might exhibit clustering patterns as different $k$-tuples reach similar alignment scores. Such clusters are clearly visible, for example, from the spectrum of multiway alignment of U.S. House in 2014 (Figure 9) at different values of $k$ ($k=2$, $k=3$, $k=4$, $k=5$, $k=16$). As the most aligned $k$-tuples are not isolated extreme occurrences, the distinguishable clusters of alignment scores suggest that members of the House might consistently align even on a wider set of issues.
Appendix H Use of the multiway_alignment Python Package
The multiway alignment Python package developed in conjunction with this research is available as open-source in GitHub. The code has $90\%$ coverage from unit tests and it is tested on Python $3.10$, $3.11$, and $3.12$. The multiway alignment package can be utilized in any Python environment using Python $\geq 3.10$ by following the steps outlined below. The package can be installed directly from the Python Package Index (PyPI) using pip from the command-line interface, by executing the following command:
\$ pip install multiway_alignment
Listing 1: Installing from pip.
Alternatively, the package can be installed by first cloning the repository containing the source code from GitHub and then installing it locally in a chosen directory, with the following commands:
\$ git clone https://github.com/letiziaia/multiway-alignment.git
\$ cd multiway-alignment
\$ pip install .
Listing 2: Installing from source.
Once the package is installed, it can be imported in a Python script or Jupyter environment to perform multiway alignment analysis. The documentation and example usage of the functions in the package is available in GitHub. Below is an example of how to import the package and use it to compute 3-way alignment between topics A, B, and C:
1import pandas as pd
2# Importing the package
3import multiway_alignment.consensus as mac
4import multiway_alignment.score as mas
5
6# Load the opinion labels for each issue to a pandas DataFrame
7df = pd.DataFrame(
8 {
9 # on topic A, the first two individuals
10 # have opinion 0, while the following two
11 # have opinion 1
12 "A": [0, 0, 1, 1],
13 "B": [0, 1, 0, 1],
14 "C": [1, 0, 1, 0],
15 }
16)
17
18# get the labels for the consensus partition over topics A, B, and C
19partition_labels = mac.get_consensus_labels(opinions=df)
20
21# compute 3-way net alignment score over topics A, B, and C
22# using adjusted mutual information without the null model
23a = mas.multiway_alignment_score(
24 df, which_score="ami", adjusted=False,
25)
Listing 3: Example Python Code.
Appendix I Alternative Formulation Using Full Consensus Partitions
The multiway alignment defined in Equation 2 compares the opinion partition defined for each topic with the consensus partition defined by the remaining topics. However, given the topics $T_{1},\ldots,T_{k}$, we could have defined multiway alignment utilizing normalized mutual information (NMI) as
$A^{*}=\frac{1}{k}\sum_{i=1}^{k}\text{NMI}(C(T_{1},\ldots,T_{k}),T_{i}),$ (6)
where NMI is the pairwise normalized mutual information (Vinh et al., 2009), $C(T_{1},\ldots,T_{k})$ is the full consensus partition of individuals defined by all the topics $T_{1},\ldots,T_{k}$, and each $T_{i}$ defines the opinion groups for the $i$-th individual topic.
Mutual information (Vinh et al., 2009) is defined in terms of entropy $H$ as
[…]
which is symmetric with respect to $X$ and $Y$, and equivalent to
[…]
By definition, any full consensus partition C perfectly satisfies homogeneity (Rosenberg and Hirschberg, 2007) with respect to each individual topic $T_{i}$, since each of the consensus groups only contains members of the same opinion group on the given topic. In terms of conditional entropy, this means that $H(T_{i}|\text{C})=0$ $\forall i$. Therefore, for each topic, the contribution of the mutual information in Equation 6 simplifies to
$\text{I}(\text{C},T_{i})=H(T_{i}),$
which can be normalized with the arithmetic average of the entropies:
$\text{NMI}(\text{C},T_{i})=\frac{2H(T_{i})}{H(T_{i})+H(\text{C})}=\frac{2}{1+\frac{H(\text{C})}{H(T_{i})}}.$
Since by definition of consensus partition, $H(\text{C})=H(T_{1},\dots,T_{k})$,
$\text{NMI}(\text{C},T_{i})=\frac{2H(T_{i})}{H(T_{i})+H(\text{C})}=\frac{2}{1+\frac{H(T_{1},\dots,T_{k})}{H(T_{i})}},$
which turns Equation 6 into:
$A^{*}=\frac{1}{k}\sum_{i=1}^{k}\frac{2}{1+\frac{H(T_{1},\dots,T_{k})}{H(T_{i})}}.$
This equivalent formulation highlights the interpretation of $A^{*}$ as a normalized comparison between the joint entropy of the variables and the individual entropies. Specifically, the term $\frac{H(T_{1},\dots,T_{k})}{H(T_{i})}\geq 1$ for any $T_{i}$. In the extreme case of perfectly overlapping opinion distributions across all the $T_{i}$, the joint entropy $H(T_{1},\dots,T_{k})$ equals the entropy of each one of the $T_{i}$, thus reducing the term $\frac{H(T_{1},\dots,T_{k})}{H(T_{i})}$ to $1$, which results in $A^{*}=1$.
Another interpretation of the normalized mutual information with respect to the consensus partition derives from the homogeneity property. By the definition of mutual information, it follows that
$H(T_{i})+H(\text{C})=H(T_{i})+H(T_{i},\text{C})$
and therefore, by the definition of conditional entropy,
$H(T_{i})+H(\text{C})=2H(T_{i})+H(\text{C}|T_{i}),$
the formula above can be also expressed in terms of the conditional entropy as
$\text{NMI}(\text{C},T_{i})=\frac{1}{1+\frac{1}{2}\frac{H(\text{C}|T_{i})}{H(T_{i})}}.$
In particular, $H(\text{C}|T_{i})=0$ (completeness property (Rosenberg and Hirschberg, 2007)) if and only if the consensus partition does not further disgregate the opinion groups formed at the topic $T_{i}$. In practice, this means that all of the individuals that belong to the same opinion group on topic $T_{i}$ are also found in the same consensus group defined by the set of topics. Since the multiway metric we defined in Equation 6 is the arithmetic average of $\text{NMI}(T_{i},\text{C})$ over a set of topics, this proves that $A^{*}=1$ if and only if the $k$ topics define the exact same partitioning of the individuals.
Finally, variation of information (Meilă, 2003) is defined in terms of entropy $H$ as
$\text{VI}(X,Y)=H(X|Y)+H(Y|X),$
which, by the homogeneity property, simplifies to
$\text{VI}(T_{i},\text{C})=H(\text{C}|T_{i}),$
showing that Equation 6 can also be expressed in terms of the variation of information by using terms of the form
$\text{NMI}(\text{C},T_{i})=\frac{1}{1+\frac{1}{2}\frac{\text{VI}(T_{i},\text{C})}{H(T_{i})}}.$
This allows to bound the terms, since variation of information is non-negative and equal to 0 when the two partitions $T_{i}$ and $C$ being compared are identical, which leads to a contribution $\text{NMI}(\text{C},T_{i})=1$. On the other hand, $\text{VI}(T_{i},\text{C})$ is upper-bounded in terms of the number of opinion groups $|C|$, or in terms of the size $N$ of the population in case $|C|=N$, which happens if the consensus partition reduces to a collection of singletons. Therefore
$0\leq\frac{\text{VI}(T_{i},\text{C})}{H(T_{i})}\leq\frac{\log(|C|)}{H(T_{i})}$
implies that
$0\leq\text{NMI}(\text{C},T_{i})\leq\frac{{H(T_{i})}}{{H(T_{i})}+\frac{1}{2}\log(|C|)}\leq 1$
I.1 Expected Alignment in Case of Equal-Sized Opinion Groups
As shown above,
$A^{*}=\frac{2}{k}\sum_{i=1}^{k}\frac{1}{1+\frac{H(\text{C})}{H(T_{i})}},$
where the entropy $H=-\Sigma_{x}p_{x}\log{p_{x}}$ of a topic is 0 if all individuals belong to the same opinion group $x$. On the other hand, the highest entropy is reached when all the opinion groups are equal-sized, as in that case the probability of belonging to one group is simply $\frac{1}{G}$, where $G$ is the number of opinion groups.
Assume a social system of $N$ individuals, and $k$ topical issues, each one allowing for the same number of equally-likely opinion groups $G$. Then, the expected $k$-wise alignment can be derived in closed form and simplifies as:
$\mathbb{E}[A^{*}]=\frac{2}{k}\sum_{i=1}^{k}\frac{1}{1+\frac{\mathbb{E}[H(\text{C})]}{H(T_{i})}}$
$=\frac{2}{k}\sum_{i=1}^{k}\frac{1}{1+\frac{\mathbb{E}[H(\text{C})]}{G(-\frac{1}{G}\log{\frac{1}{G}})}}$
$=\frac{2}{1+k}.$